When resistors 1 and 2 are connected in series, the equivalent resistance is 14.3 Ω. When they are connected in parallel, the equivalent resistance is 2.70 Ω. What are (a) the smaller resistance and (b) the larger resistance of these two resistors?
Answer #1

Let the resistors be R1 and R2

Equivalent resistance in series is given as

R1 + R2 = 14.3 eq-1

R2 = 14.3 - R1                                               eq-2

Equivalent resistance in parallel is given as

R1 R2/(R1 + R2) = 2.70

Using eq-1

R1 R2/(14.3) = 2.70

R1 R2 = 38.61

Using eq-2

R1 (14.3 - R1) = 38.61

R1 = 3.61 ohm

b)

Using eq-2

R2 = 14.3 - R1

R2 = 14.3 - 3.61

R2 = 10.7 ohm